# Martian World Models: Controllable Video Synthesis with Physically Accurate 3D Reconstructions

### v1

@@ 5 Experiments
Martian Terrain Video Generation. We quantitatively evaluate the performance of MarsGen, our model fine-tuned on the MarsSynthSim dataset, with a focus on both visual fidelity and 3D geometric consistency. We assess visual quality using FID (Fréchet Inception Distance) [18] and FVD (Fréchet Video Distance) [54], and evaluate geometric consistency through the 2D Warp Error and PSNR. A key capability of MarsGen is its precise camera pose control, which enables us to directly evaluate its novel view synthesis (NVS) performance. Table 1 compares MarsGen against state-of-the-art image-to-video and camera-controlled video generation models. MarsGen consistently outperforms all baselines in terms of visual fidelity, achieving the lowest FID and FVD. In terms of geometric consistency, MarsGen achieves the lowest 2D Warp Error and competitive PSNR, indicating superior preservation of 3D structure during generation. For novel view synthesis, MarsGen also shows superior performance across all metrics, demonstrating the benefit of 3D-aware training on MarsSynthSim. For a qualitative comparison, please see Appendix B. In contrast, general-purpose video models often struggle to maintain consistent spatial geometry under camera motion due to the absence of explicit 3D supervision.
Figure 5: Qualitative comparison of point cloud reconstruction from a Martian input view. Our M3arsSynth engine (top right) produces a coherent point cloud accurately capturing terrain. In contrast, the VGGT [58] model (bottom right) exhibits significant misalignment and artifacts.
Geometry Initialization Methods. We evaluate alternative approaches for initializing scene geometry within our M3arsSynth pipeline (used to generate the MarsSynthSim dataset), comparing traditional Structure-from-Motion (SfM) pipelines (e.g., COLMAP) and recent transformer-based methods (e.g., MASt3R) against our metric-aware initialization. Our method combines pre-trained vision foundation models to robustly extract camera and depth priors from challenging Martian stereo imagery. As indicated in Table 2, our pipeline achieves 100% data utilization and reconstructs dense point clouds while maintaining competitive reprojection accuracy and a reasonable runtime. In contrast, COLMAP experiences partial frame failures, and MASt3R, despite generating dense reconstructions, exhibits significant reprojection errors, potentially due to overfitting to unreliable depth priors. Furthermore, we observed that point clouds derived directly from the VGGT model often contain significant artifacts (see Figure 5). We attribute these artifacts to out-of-distribution challenges encountered by the model. Consequently, our reconstruction pipeline utilizes VGGT for intrinsic estimation rather than for direct point cloud export for 3D-GS initialization.
@@ A.3 Details of M3arsSynth construction
$\mathcal{L}_{\text{depth}}=||D_{render}-D^{*}_{\text{Metric Depth}}||_{1}$ (7)
@@
Camera Trajectory Synthesis and Motion Description   A cornerstone of generating diverse and informative video sequences for the M3arsSynth dataset is the meticulous design and execution of virtual camera trajectories. We begin by defining a repertoire of canonical camera trajectory types. These foundational trajectories, mathematically represented as a sequence of 6-Degrees-of-Freedom (6-DOF) poses $\mathcal{M}_{\text{traj}}=\{(R_{t},T_{t})\in\text{SE(3)}\mid t=1,\dots,N\}$, where $R_{t}$ signifies the camera’s rotation matrix and $T_{t}$ denotes its translation vector at each discrete timestamp $t$, are engineered to encompass a wide spectrum of motion profiles. The spatial extent, or the overall scale and reach, of the predefined canonical trajectories is not fixed; instead, it is dynamically adjusted in response to the specific geometric characteristics of each individual 3D reconstructed Martian scene. This adaptation is primarily driven by the depth information derived from the reconstructed 3D model. Specifically, for regions within a scene that are identified as being in the near-field (characterized by relatively smaller depth values from the camera’s perspective), the corresponding segments of the canonical trajectories are programmatically contracted or scaled down. Conversely, for regions designated as far-field (characterized by significantly larger depth values, indicating distant terrain elements or horizons), the trajectory segments are expanded or scaled up. This depth-adaptive scaling strategy is paramount for ensuring that the synthesized video data effectively and consistently covers the scene’s content at appropriate levels of detail. Following the generation of these adaptively scaled trajectories, the precise 6-DOF pose parameters for each frame serve as the quantitative foundation from which natural language descriptions detailing the camera’s motion characteristics are subsequently derived, forming a key component of the textual modality within the M3arsSynth dataset.
@@
Scene Content Captioning   The M3arsSynth dataset incorporates rich textual descriptions to enable and enhance multimodal learning. While depth and normal maps are directly derived from the 3D reconstructed scenes, and camera motion characteristics are derived from the 6-DOF pose parameters of the trajectories, the acquisition of descriptive scene content captions involves a sophisticated process leveraging a Vision Language Model (VLM). We generate scene content captions by applying a VLM, referenced as ChatGPT-4o, to selected views from the synthesized video sequences. The generation process is guided by specific system prompts provided to the VLM. These prompts include the input image, the classification of the Martian terrain depicted (e.g., "Regolith/Rocky Terrain", "Dunes/Ripples (Sand/Dust)", as shown in Figure 4 of the paper), and a basic descriptive outline of the scene. By conditioning the VLM with this structured input, we obtain detailed and contextually relevant textual descriptions of the visual content. This methodology ensures that the textual modality is not only accurate but also aligned with the visual and geometric data, thereby creating a cohesive multimodal dataset suitable for training models like MarsGen for controllable video synthesis.
@@ Appendix B Additional Experimental Results and Analysis
Qualitative Comparisons of Video Generation.   The main paper presented quantitative comparisons of our generator against image-to-video and camera-controlled image-to-video models. This appendix provides additional quantitative results against other video generation models. Sora and ViewCrafter, for instance, evidently lack specialized modeling for dynamic Martian scenes, leading to uncontrollable video sequences inconsistent with the theme. This further validates the significance of our proposed dataset.
@@
Qualitative Comparisons of 3D Reconstruction Pipelines.   Fig. B illustrates a qualitative comparison of 3D reconstruction outputs from different methodologies when applied to demanding Martian stereo image pairs. The top row of the figure presents results achieved by our M3arsSynth pipeline, which consistently demonstrates the ability to generate coherent and detailed 3D reconstructions of the Martian terrain. These outputs effectively capture the complex geometry and features of the landscape.

### v2

@@ 5 Experiments
Martian Terrain Video Generation. We quantitatively evaluate the performance of MarsGen, our model fine-tuned on the MarsSynthSim dataset, with a focus on both visual fidelity and 3D geometric consistency. We assess visual quality using FID (Fréchet Inception Distance) heusel2017gans and FVD (Fréchet Video Distance) unterthiner2018towards, and evaluate geometric consistency through the 2D Warp Error and PSNR. A key capability of MarsGen is its precise camera pose control, which enables us to directly evaluate its novel view synthesis (NVS) performance. Table 1 compares MarsGen against state-of-the-art image-to-video and camera-controlled video generation models. MarsGen consistently outperforms all baselines in terms of visual fidelity, achieving the lowest FID and FVD. In terms of geometric consistency, MarsGen achieves the lowest 2D Warp Error and competitive PSNR, indicating superior preservation of 3D structure during generation. For novel view synthesis, MarsGen also shows superior performance across all metrics, demonstrating the benefit of 3D-aware training on MarsSynthSim. For a qualitative comparison, please see Appendix B. In contrast, general-purpose video models often struggle to maintain consistent spatial geometry under camera motion due to the absence of explicit 3D supervision.
Figure 5: Qualitative comparison of point cloud reconstruction from a Martian input view. Our M3arsSynth engine (top right) produces a coherent point cloud accurately capturing terrain. In contrast, the VGGT wang2025vggt model (bottom right) exhibits significant misalignment and artifacts.
Geometry Initialization Methods. We evaluate alternative approaches for initializing scene geometry within our M3arsSynth pipeline (used to generate the MarsSynthSim dataset), comparing traditional Structure-from-Motion (SfM) pipelines (e.g., COLMAP) and recent transformer-based methods (e.g., MASt3R) against our metric-aware initialization. We use the 2D reprojection error which measures the pixel distance between an observed point in the original image and its corresponding 3D point when re-projected back onto the image using the estimated camera parameters. It thereby jointly assesses the geometric accuracy of the 3D model and the estimated camera pose. Our method combines pre-trained vision foundation models to robustly extract camera and depth priors from challenging Martian stereo imagery. As indicated in Table 2, our pipeline achieves 100% data utilization and reconstructs dense point clouds while maintaining competitive reprojection accuracy and a reasonable runtime. In contrast, COLMAP experiences partial frame failures, and MASt3R, despite generating dense reconstructions, exhibits significant reprojection errors, potentially due to overfitting to unreliable depth priors. Furthermore, we observed that point clouds derived directly from the VGGT model often contain significant artifacts (see Figure 5). We attribute these artifacts to out-of-distribution challenges encountered by the model. Consequently, our reconstruction pipeline utilizes VGGT for intrinsic estimation rather than for direct point cloud export for 3D-GS initialization.
@@ Acknowledgements
This research has been supported by computing support on the Vista GPU Cluster through the Center for Generative AI (CGAI) and the Texas Advanced Computing Center (TACC) at the University of Texas at Austin.
@@ A.3 Details of M3arsSynth construction
[…]
@@
Camera Trajectory Synthesis and Motion Description   A cornerstone of generating diverse and informative video sequences for the M3arsSynth dataset is the meticulous design and execution of virtual camera trajectories. We begin by defining a repertoire of canonical camera trajectory types. These foundational trajectories, mathematically represented as a sequence of 6-Degrees-of-Freedom (6-DOF) poses $\mathcal{M}_{\text{traj}}=\{(R_{t},T_{t})\in\text{SE(3)}\mid t=1,\dots,N\}$, where $R_{t}$ signifies the camera’s rotation matrix and $T_{t}$ denotes its translation vector at each discrete timestamp $t$, are engineered to encompass a wide spectrum of motion profiles. Our trajectory generation process begins by using the two camera poses, solved from the original stereo pair, as start and end keyframes. We then synthesize a variety of smooth camera paths between these keyframes by applying predefined interpolation methods that follow canonical motion profiles. The spatial extent, or the overall scale and reach, of the predefined canonical trajectories is not fixed; instead, it is dynamically adjusted in response to the specific geometric characteristics of each individual 3D reconstructed Martian scene. This adaptation is primarily driven by the depth information derived from the reconstructed 3D model. Specifically, for regions within a scene that are identified as being in the near-field (characterized by relatively smaller depth values from the camera’s perspective), the corresponding segments of the canonical trajectories are programmatically contracted or scaled down. Conversely, for regions designated as far-field (characterized by significantly larger depth values, indicating distant terrain elements or horizons), the trajectory segments are expanded or scaled up. This depth-adaptive scaling strategy is paramount for ensuring that the synthesized video data effectively and consistently covers the scene’s content at appropriate levels of detail. Following the generation of these adaptively scaled trajectories, the precise 6-DOF pose parameters for each frame serve as the quantitative foundation from which natural language descriptions detailing the camera’s motion characteristics are subsequently derived, forming a key component of the textual modality within the M3arsSynth dataset.
@@
Scene Content Captioning   The M3arsSynth dataset incorporates rich textual descriptions to enable and enhance multimodal learning. While depth and normal maps are directly derived from the 3D reconstructed scenes, and camera motion characteristics are derived from the 6-DOF pose parameters of the trajectories, the acquisition of descriptive scene content captions involves a sophisticated process leveraging a Vision Language Model (VLM). We generate scene content captions by applying a VLM, referenced as ChatGPT-4o, to selected views from the synthesized video sequences. Guided by expertise in Martian exploration from our authors, we developed structured prompts for the VLM to ensure high-quality annotations. These prompts include the input image, the classification of the Martian terrain depicted (e.g., "Regolith/Rocky Terrain", "Dunes/Ripples (Sand/Dust)", as shown in Figure 4 of the paper), and a basic descriptive outline of the scene. By conditioning the VLM with this structured input, we obtain detailed and contextually relevant textual descriptions of the visual content. We manually validated a random sample of 20% of the annotations and found no significant errors, which attests to the robustness of our pipeline.This methodology ensures that the textual modality is not only accurate but also aligned with the visual and geometric data, thereby creating a cohesive multimodal dataset suitable for training models like MarsGen for controllable video synthesis.
@@ Appendix B Additional Experimental Results and Analysis
B.1 Qualitative Comparisons of Video Generation
The main paper presented quantitative comparisons of our generator against image-to-video and camera-controlled image-to-video models. This appendix provides additional quantitative results against other video generation models. Sora and ViewCrafter, for instance, evidently lack specialized modeling for dynamic Martian scenes, leading to uncontrollable video sequences inconsistent with the theme. This further validates the significance of our proposed dataset.
@@ B.1 Qualitative Comparisons of Video Generation
B.2 Qualitative Comparisons of 3D Reconstruction Pipelines
Fig. B illustrates a qualitative comparison of 3D reconstruction outputs from different methodologies when applied to demanding Martian stereo image pairs. The top row of the figure presents results achieved by our M3arsSynth pipeline, which consistently demonstrates the ability to generate coherent and detailed 3D reconstructions of the Martian terrain. These outputs effectively capture the complex geometry and features of the landscape.
@@ B.2 Qualitative Comparisons of 3D Reconstruction Pipelines
B.3 Discussion on Geometric Initialization Component Choices
Our experiments, summarized in Table A, validate our component choices for geometric initialization. Our primary goal is accurate, metric-scale reconstruction, which is critical for Martian terrain analysis but not a direct output of methods like VGGT which predicts normalized depth.
The ablation study highlights the limitations of alternatives. The full MASt3R(full) pipeline and its intrinsic estimation (Ours(MASt3R intrin.)) are unstable (46.980 px and 2.016 px, respectively) due to inaccurate 2D correspondences. In contrast, we found VGGT provides more stable intrinsic estimations, likely due to its DINOv2-initialized encoder and normalized training targets.
Furthermore, using VGGT’s outputs directly is suboptimal. Scaling its normalized depth to metric scale (VGGT(metric depth)) is also suboptimal (1.748 px). Using its predicted extrinsics (Ours(VGGT extrin.)) also yields poor results (2.113 px). This is because these non-metric scale poses are fundamentally incompatible with the metric depth priors from Metric3D v2, which are essential to our pipeline. Our approach successfully integrates the strengths of these models while ensuring metric-scale consistency.
Table A: Ablation study on geometric initialization components. Our full method achieves the lowest 2D reprojection error, indicating the most accurate geometric setup.
MASt3R(full) VGGT(metric depth) Ours(MASt3R intrin.) Ours(VGGT extrin.) Ours
2D Reproj. $\downarrow$ 46.980 1.748 2.016 2.113 0.770
B.4 Discussion on Generalization to Lunar and Terrestrial Environments
While our primary motivation stems from Martian exploration, our method is designed for planetary scenes with sparse views and low-texture surfaces. We evaluated our pipeline on similar environments, including public lunar data, performing novel view synthesis from two views.
Furthermore, to test robustness in a general sparse-view context, we performed novel view synthesis on the terrestrial RealEstate10K dataset using only two-view inputs. We benchmarked our approach against Splatt3R and InstantSplat.
The preliminary results, shown in Table B, demonstrate our method’s effectiveness in these challenging sparse-view conditions on both lunar and terrestrial data.
Table B: Sparse-view novel view synthesis benchmarks on Lunar and RealEstate10K data.
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Lunar Ours 29.60 0.920 0.213
RealEstate10K Splatt3R 15.11 0.492 0.442
InstantSplat 19.64 0.560 0.291
Ours 20.81 0.717 0.264